OpenStudy (anonymous):

Add & Simplify 3a/a+b + 3b/a+b

OpenStudy (rulnick):

Answer is 3.

OpenStudy (rulnick):

... 3a/(a+b) + 3b/(a+b) = (3a+3b)/(a+b) = 3(a+b)/(a+b) = 3.